# Leveraging Temporal Contextualization for Video Action Recognition

### v1

@@ 5.4 Qualitative Analysis
Figure 7: Attention visualization. While ViFi-CLIP fails to attend to the movement of hands moving farther away and misinterprets the action as colliding, TC-CLIP correctly predicts with temporal consistency.
@@ 6 Conclusion
We have introduced TC-CLIP, a novel video understanding paradigm that leverages holistic video information within the encoding process. Unlike prior approaches that access only a limited range of tokens, our temporal contextualization summarizes informative tokens from the entire video and utilizes them during the attention operation. While these tokens are exploited to infuse temporal information on the vision side, they also serve as a source for video-conditional text prompting, enhancing instance-wise context on the text side. Extensive experiments and analysis on diverse video recognition benchmarks and evaluation protocols demonstrate the superiority of TC-CLIP and its design choices.
Limitation. Due to resource constraints, we verified TC-CLIP only using ViT-B. We will further explore its scalability with ViT-L and ViT-H.
@@ Appendix 0.A Fine-tuning with the Kinetics-400 Pretrained Model
We additionally compare TC-CLIP with the state-of-the-art methods on few-shot and base-to-novel settings using the Kinetics-400 pretrained model. In this setting, all the models are first pretrained on the Kinetics-400 dataset and subsequently fine-tuned on each dataset. Table 8 summarizes the comparison results in the few-shot action recognition regime. Notably, TC-CLIP surpasses other methods by a large margin of at least 2%p in the average. Furthermore, Table 9 shows the comparison results in the base-to-novel generalization setup. In an average over all datasets, TC-CLIP shows distinct gaps from the baseline model ViFi-CLIP [26] by 3.5%p, 6%p, and 4.8%p in base, novel, and harmonic mean (HM), respectively.
K-400 fully-supervised SSv2 16-shot
Method All Temporal All Temporal
ViFi-CLIP [26] 83.9 87.8 12.4 25.9
TC-CLIP (Ours) 85.2 ($+$1.3) 89.2 ($+$1.4) 14.0 ($+$1.6) 29.9 ($+$4.0)
Table 11: VP layer selection ablation.
We report $K$-shot training results where the top-1 accuracy in each dataset is averaged over $K=2,4,8,16$. The late-stage prompting design generalizes well.
@@
Table 12: VP initialization ablation by varying initialization cases of the VP layer and prompt vectors. We report $K$-shot training results where the top-1 accuracy is averaged over $K=2,4,8,16$. In general, it is advantageous to avoid random initialization.
@@
CLIP a photo of a 63.9 90.7 9.8 54.8
Random a photo of a 63.5 90.5 9.9 54.6
CLIP Random 62.5 90.2 9.5 54.1
Appendix 0.B Temporal Subset Analysis
We adopt the temporal subset analysis suggested in [27] to further analyze the temporal modeling ability of trained models. In this setting, the temporal subset consists of specific action classes that require more temporal information to recognize them. After shuffling frames in time and measuring whether the human annotator can still recognize the action, the classes that cannot be recognized are added to the temporal subset. The resulting temporal classes in Kinetics-400 and SSv2 are a total of 32 classes and 16 classes, respectively. Table 10 compares the top-1 accuracy using full validation splits and the temporal subsets. We use fully-supervised models for Kinetics-400 and 16-shot-trained models for SSv2. Our model shows unique gains over ViFi-CLIP [26] for the temporally challenging videos, more than when evaluated across the entire validation set. This demonstrates the superiority of TC-CLIP in handling temporal information.
Appendix 0.C More Ablation Study on VP
@@
VP layer selection. Table 0.A presents the ablation study on the layer selection in the VP module. We compare the few-shot classification accuracy by varying the layer indices of the text and vision inputs $\{L_{\text{text}},L_{\text{vision}}\}$ in the VP module $f_{\theta_{\text{VP}}}(\mathbf{p}^{L_{\text{text}}-1},\mathbf{s}^{L_{\text{vision}}}_{\text{proj}})$. We observe that applying conditional prompting at the early stage ($L_{\text{text}}=1$) does not generalize well, regardless of the vision layer index. This suggests the early-stage prompting design is hard to generalize in a full fine-tuning scenario, possibly because CLIP was initially trained in a vision-text late-alignment fashion. The middle- and late-stage prompting designs are also compared, and the last layer indices are finally chosen for both modalities.
@@
VP initialization. Table 0.A presents the ablation study on the initialization strategy of the VP layer and learnable prompt vectors. By default, we initialize the weight of the VP module using the weight from the last layer of the CLIP text encoder since randomly initializing it occasionally incurs unstable training results in the few-shot scenario. Similarly, it is beneficial to initialize the learnable prompt vectors using the prompt template “a photo of a” as generally done in the prompt tuning methods [37, 17].
Appendix 0.D Impact of Weight-space Ensemble
@@
In Fig. 8, we evaluate the effectiveness of weight ensembling by varying the ensemble ratio $w$ from 0 to 1 with a step size of 0.1. Specifically, the backbone weights of both vision and text encoders are linearly interpolated between CLIP and fine-tuned model, i.e., $\theta_{w}=(1-w)\cdot\theta_{\text{CLIP}}+w\cdot\theta_{\text{fine-tuned}}$. The $y$-axis displays the average accuracy on the zero-shot video datasets and the $x$-axis displays the accuracy on the fine-tuning dataset K-400. Our TC-CLIP achieves a better trade-off than the baseline as our curve is always on top of the baseline’s curve. This demonstrates that our model takes more advantages from weight ensembling. We choose $w=0.7$ as our final ensemble ratio.
@@
Figure 10: Attention visualization of TC-CLIP in comparison with ViFi-CLIP [26] on Kinetics-400, Kinetics-600, and SSv2 datasets. (a)–(b): TC-CLIP tends to focus more on fast-moving parts such as hands and arms. (c)–(d): While ViFi-CLIP dominantly attends to the most salient regions, TC-CLIP attends to multiple objects based on inter-object relationships relevant to the occurring actions. (e)–(f): TC-CLIP consistently attends to the main object with deformations throughout the video.
Appendix 0.E More Visualizations of Context Tokens and Attentions
Context token visualization. Fig. 9 visualizes the seed tokens and context tokens from the last layer of the vision encoder in TC-CLIP. As can be seen, the seed tokens mainly consist of patch tokens from the most informative regions in each frame, which contain the foreground, such as a person, animals, hands, and objects. To visualize each context token, we colorize its corresponding source token positions using the average color of the input image patches of that region. It is noteworthy that a single context token (highlighted in red) successfully tracks and summarizes a specific object or part throughout the entire video.
Attention visualization. Fig. 10 visualizes the attention maps of TC-CLIP compared to ViFi-CLIP [26]. As shown in Fig. 10(a)–(b), TC-CLIP tends to focus more on dynamically moving parts, such as hands and arms, during the action of throwing or shooting objects. Furthermore, TC-CLIP attends to multiple objects simultaneously based on inter-object relationships, as illustrated in Fig. 10(c)–(d). For example, during actions like “swinging baseball bat,” it effectively attends to both the bat and the baseball being struck. This highlights the TC-CLIP’s capability to understand complex interactions based on a temporal context in dynamic scenes, which is potentially more helpful than ViFi-CLIP, which only attends to the most salient region in each frame. Finally, Fig. 10(e)–(f) shows TC-CLIP’s consistent attention towards objects with deformations across video frames, which are more striking than ViFi-CLIP’s.
Appendix 0.F Experimental Setup Details
0.F.1 Dataset Details
@@
0.F.2 Implementation Details
During the bipartite matching, we start with the seed tokens arranged based on the [CLS] token attention values in each frame. These tokens are then divided into two sets by alternating positions. Subsequently, $r$ pairs of tokens with the highest cosine similarity are merged by averaging their features, and remaining two sets are then concatenated back together. This process is repeated iteratively, employing a constant $r$ scheduling for every iteration with an exception in the final iteration to ensure that the number of final context tokens becomes $k$. Following [37, 17], the learnable prompts are initialized with the prompt “a photo of a”. For training recipes, we follow [26] for zero-shot, few-shot, and fully-supervised settings and follow [9] for base-to-novel generalization. By default, we use the AdamW optimizer with momentum betas of (0.9, 0.98) and a weight decay of 0.001. The initial learning rate of the VP module is 10×10$\times$10 × larger than the base learning rate in each setting. Training configurations and evaluation metrics in each protocol are specified below.
@@ 0.F.2 Implementation Details
Few-shot action recognition. We adopt the $K$-shot training splits from [26] that randomly samples $K=2,4,8,16$ videos from each class on HMDB-51, UCF-101, and SSv2. The models are evaluated using the first validation split of HMDB-51 and UCF-101 and the full validation split of SSv2. The base learning rate is set to $2\times 10^{-6}$ and is decayed to $2\times 10^{-8}$. The batch size is 64, and the total number of epochs is set to 50, starting with 5 linear warmup epochs.

### v2

@@ 3.2 Analysis and Discussion
Figure 7: Attention visualization. While ViFi-CLIP fails to attend to the hands moving away and misinterprets the action as colliding, TC-CLIP correctly predicts by exploiting temporal consistency.
Context token visualization. Fig. 6 visualizes the seed tokens and context tokens from the last layer of the vision encoder in TC-CLIP. In this video, the informative regions regarding the action of disc golfing in each frame, including the person and the disc, are selected as seed tokens. To visualize each context token, we colorize its corresponding source token positions using the average color of the input video patches of that region. Note that a single context token (highlighted in red) successfully tracks the disc across multiple frames.
Attention visualization. Fig. 7 visualizes the attention map of ViFi-CLIP [35] and TC-CLIP on the SSv2 dataset. In this video, where two hands grab objects and then move away, ViFi-CLIP [35] fails to attend to the hands from the middle of the sequence and misinterprets the action as colliding with each other. In contrast, TC-CLIP considers the temporal context across the sequence by its design, and thus consistently attends to the hands throughout the entire video and correctly predicts the action as moving away from each other.
4 Related Work
Token aggregation. Recent studies on token aggregation aim to reduce the number of tokens given to image Transformers [34, 21, 31, 24, 48, 49, 9, 27, 29, 51, 1, 20] and video Transformers [37, 40, 36, 23, 7, 6] for efficient inference. While some of these approaches train additional networks for token selection [34, 21, 40] or fusion [31, 37], we focus on parameter-free approaches, categorized into token pruning and merging. Pruning-based methods [24, 48, 49, 9] eliminate uninformative tokens by measuring their importance using a metric such as a self-attention score, whereas merging-based methods combine tokens with large semantic similarity into single units using clustering algorithms such as $k$-means [29], DPC-KNN [51], and bipartite soft matching [1, 36, 23]. To minimize information loss, several studies [27, 7, 6] consider both token importance and similarities as aggregation criteria. While our primary goal is not to improve efficiency, we employ both pruning and merging techniques to connect relevant tokens and summarize essential contexts within videos. Although there are semantic segmentation studies [46, 47, 28] that link relevant spatial data, they rely on learnable cluster centers with slot- or cross-attention blocks, and thus differ from our approach.
Prompt learning. Several studies on prompt learning [13, 53, 52, 18, 35, 43, 15] transfer VLMs to downstream tasks by optimizing a discrete set of prompt vectors. In video recognition, [15] has introduced text prompt tuning, while ViFi-CLIP [35] and Vita-CLIP [43] perform prompting in both vision and text branches. However, these prompt vectors are separately optimized and not shared across the modalities. In image recognition, Co-CoOp [52] performs an instance-conditional prompt tuning by explicitly conditioning text prompts on the [CLS]
tokens from image instances. MaPLe [18] learns multi-modal prompting by sharing layerwise context prompts for both branches. Unlikely, we generate video-conditional prompts by utilizing contextualized tokens as vision inputs and injecting summarized video information into text prompt vectors.
@@ 5 Conclusion
We have introduced TC-CLIP, a novel video understanding paradigm that leverages holistic video information within the encoding process. Unlike prior approaches that access only a limited range of tokens, the proposed temporal contextualization summarizes informative tokens from the entire video and utilizes them for attention operations. While these tokens are employed to infuse temporal information on the vision side, they also serve as a source for video-conditional text prompts, thus enhancing the instance-wise context on the text side. Extensive experiments and analyses on diverse benchmarks and evaluation protocols demonstrate the superiority of TC-CLIP and justify its design choices.
Acknowledgements
Experiments are based on the NAVER Smart Machine Learning NSML [19] platform. This research was partly supported by the Bio & Medical Technology Development Program of the National Research Foundation (NRF) (No. 2021M3A9E4080782) and the IITP grants [No. RS-2021-II212068; No. RS-2021-II211343] funded by the Korean government (MSIT).
@@ Appendix 0.A Fine-tuning with the Kinetics-400 Pretrained Model
Tables 11 and 12 present the comparison results using the K-400 pretrained model on the few-shot and base-to-novel settings. All the models are first pretrained on the Kinetics-400 dataset and subsequently fine-tuned on each dataset. TC-CLIP demonstrates superior performance over all other methods by significant margins. Particularly in the base-to-novel setup, TC-CLIP outperforms ViFi-CLIP [35] with notable gaps of 3.5%p, 6%p, and 4.8%p in the base, novel, and harmonic mean (HM) on average, respectively.
Appendix 0.B More Ablation Study on VP
Table 13: Video-conditional Prompting (VP) ablation. We report $K$-shot training results where the top-1 accuracy in each dataset is averaged over $K=2,4,8,16$. Default settings are marked in gray.
(a) Number of prompt vectors.
Case HMDB UCF SSv2 All
2 62.0 90.1 9.4 53.8
4 63.9 90.7 9.8 54.8
8 63.7 90.4 9.8 54.7
(b) Vision input token selection.
Case HMDB UCF SSv2 All
CLS tokens from all frames 63.3 90.3 9.8 54.4
GAP tokens from all frames 63.7 90.5 9.8 54.7
Context tokens 63.9 90.7 9.8 54.8
(c) Input layer selection.
@@ Appendix 0.B More Ablation Study on VP
(d) Layer and prompt initialization.
@@
CLIP weight a photo of a 63.9 90.7 9.8 54.8
Random init. a photo of a 63.5 90.5 9.9 54.6
CLIP weight Random init. 62.5 90.2 9.5 54.1
Table 14: Computational cost–performance trade-off of VP design. In the case of vision-text late-fusion design, class name embeddings are pre-computed. Models are evaluated on the zero-shot setting without the weight ensemble. Costs are measured using a single A6000 GPU.
K-600 GFLOPs Latency (s)
Case Top-1 Vision Text Cross-modal Training Inference
Vision-text late-fusion 70.9 301 2.96 0.06 0.54 (1.00$\times$) 0.042 (1.00$\times$)
Video-conditional Prompting (VP) 72.7 (+1.8) 301 2.96 0.06 0.58 (1.07$\times$) 0.045 (1.07$\times$)
Table 13 examines the design choice of VP in TC-CLIP on the few-shot setting.
Vision token selection. Using context tokens in VP yields better results than employing [CLS] tokens or global average pooled (GAP) tokens from all frames. This demonstrates that proper contextualization of vision features is essential to transfer the video information to the text side.
@@
Input layer selection. We vary the layer indices of the text and vision inputs $\{L_{\text{text}},L_{\text{vision}}\}$ in the VP module $f_{\theta_{\text{VP}}}(\mathbf{p}^{L_{\text{text}}-1},\mathbf{s}^{L_{\text{vision}}}_{\text{proj}})$. We observe that conditional prompting at the early stage ($L_{\text{text}}=1$) does not generalize well, regardless of the vision layer index. The early-stage prompting design is hard to generalize in a full fine-tuning scenario, possibly because CLIP was initially trained in a vision-text late-alignment fashion. Consequently, we choose the late-stage prompting by adopting the last layers for both modalities.
@@
Layer and prompt initialization. We initialize the VP module’s weight using the weight from the last layer of the CLIP text encoder because random initialization often results in unstable training results in the few-shot scenario. Similarly, it is beneficial to initialize the learnable prompt vectors using the prompt template “a photo of a” following several prompt tuning methods [52, 18].
Computational cost analysis. Although VP requires the instance-conditional computation of text embeddings, the added cost is minor. As in Table 14, for a pair of video and text inputs, the GFLOPs required by the text encoder cost only about 1% of those needed by the vision encoder. Given these minimal text-related costs, VP adds only an extra 0.07$\times$ in latency compared to the vision-text late-fusion design using pre-computed text embeddings. Considering the observed performance gain, this is an acceptable trade-off.
Table 15: Comparison with state-of-the-arts on zero-shot action recognition using ViT-L/14. All the models are trained on Kinetics-400 and directly evaluated on other datasets. † denotes that the results are reproduced with our implementation. The best results are in bold-faced numbers, and the second-best ones are underlined.
Method WE HMDB-51 UCF-101 K600 (Top-1) K600 (Top-5) All (Top-1)
ViFi-CLIP [35] 55.6 $\pm$ 0.5 86.1 $\pm$ 0.8 77.8 $\pm$ 0.9 95.6 $\pm$ 0.2 73.2
TC-CLIP (Ours) 56.1 $\pm$ 0.3 86.9 $\pm$ 0.9 80.1 $\pm$ 0.7 96.5 $\pm$ 0.1 74.4
ViFi-CLIP [35] 55.8 $\pm$ 0.7 88.1 $\pm$ 1.3 81.1 $\pm$ 0.7 96.7 $\pm$ 0.1 75.0
Open-VCLIP [44] 59.0 $\pm$ 0.6 87.6 $\pm$ 1.2 81.1 $\pm$ 0.8 96.3 $\pm$ 0.3 75.9
TC-CLIP (Ours) 57.1 $\pm$ 0.7 88.9 $\pm$ 0.9 83.1 $\pm$ 0.7 97.3 $\pm$ 0.1 76.4
K-400 fully-supervised SSv2 16-shot
Method All Temporal All Temporal
ViFi-CLIP [35] 83.9 87.8 12.4 25.9
TC-CLIP (Ours) 85.2 ($+$1.3) 89.2 ($+$1.4) 14.0 ($+$1.6) 29.9 ($+$4.0)
Appendix 0.C Scalability with ViT-L/14
Table 15 shows the zero-shot performance comparison using CLIP ViT-L/14 as a backbone. In the case of using WE, our model outperforms ViFi-CLIP [35] and Open-VCLIP [44] by 1.4%p and 0.5%p on average, respectively.
Appendix 0.D Temporal Subset Analysis
We adopt the temporal subset analysis suggested in [38] to further analyze the temporal modeling ability of trained models. The temporal subset consists of several action classes that require more temporal information to recognize them, i.e., the classes of videos that cannot be recognized by human annotators after randomly shuffling the frames. As shown in Table 16, TC-CLIP’s gains over ViFi-CLIP [35] on the temporal subsets are more substantial than the gains when evaluated on the full validation splits, demonstrating the superiority in handling temporal information.
Appendix 0.E Impact of Weight-space Ensembling
@@
In Fig. 8, we evaluate the effectiveness of weight ensembling by varying the ensemble ratio $w$ from 0 to 1 with a step size of 0.1. Specifically, the backbone weights of both vision and text encoders are linearly interpolated between CLIP and fine-tuned model, i.e., $\theta_{w}=(1-w)\cdot\theta_{\text{CLIP}}+w\cdot\theta_{\text{fine-tuned}}$. The $y$-axis shows the average accuracy on the zero-shot video datasets, and the $x$-axis means the accuracy on the fine-tuning dataset K-400. Our model achieves a better trade-off than the baseline as our curve is always on top of the baseline’s curve. This demonstrates that our model takes more advantages from weight ensembling. We choose $w=0.7$ as our final ensemble ratio.
@@
Figure 10: Attention visualization of TC-CLIP in comparison with ViFi-CLIP [35] on Kinetics-400, Kinetics-600, and SSv2 datasets using [CLS] token as a query in each frame. (a)–(b): TC-CLIP tends to focus more on fast-moving parts such as hands and arms. (c)–(d): While ViFi-CLIP dominantly attends to the most salient regions, TC-CLIP attends to multiple objects based on inter-object relationships relevant to the occurring actions. (e)–(f): TC-CLIP consistently attends to the main object with deformations throughout the video.
Figure 11: Attention visualization of TC-CLIP in comparison with various temporal information learning approaches on SSv2 dataset. We visualize the attention map in the last vision encoder layer using a ball (top) and a hand (bottom) as a query (denoted with red boxes). To visualize the attention map from TC, we assign attention values of context tokens to their corresponding source patch token positions. Unlike other approaches, our method successfully highlights informative regions globally over frames.
Appendix 0.F More Visualizations of Context Tokens and Attentions
Context token visualization. Fig. 9 visualizes the seed tokens and context tokens from the last layer of the vision encoder in TC-CLIP. The seed tokens mainly consist of patch tokens from the most informative regions in each frame, often corresponding to the foreground, such as a person, animals, hands, and objects. To visualize each context token, we colorize its corresponding source token positions using the average color of the input image patches of that region. It is noteworthy that a single context token (highlighted in red) successfully tracks and summarizes a specific object or part throughout the entire video.
Class token attention visualization. Fig. 10 visualizes the attention maps of TC-CLIP compared to ViFi-CLIP [35] using [CLS] token as a query in each frame. As shown in Fig. 10(a)–(b), during the action of throwing or shooting objects, TC-CLIP tends to focus more on dynamically moving parts such as hands and arms. Furthermore, as in Fig. 10(c)–(d), TC-CLIP highlights multiple objects simultaneously based on inter-object relationships. During actions like “swinging baseball bat,” TC-CLIP focuses on both the bat and the baseball being struck, whereas ViFi-CLIP only highlights salient areas in individual frames. Fig. 10(e)–(f) also shows TC-CLIP’s consistent attention towards objects with deformations across frames, which is more striking than ViFi-CLIP’s.
Patch token attention visualization. Fig. 11 shows the attention maps of TC-CLIP compared to other temporal modeling approaches [30, 43, 44] by using a patch token as a query. To visualize the attention map from TC, we assign attention values of context tokens to their corresponding source patch token positions. In both examples, the token interactions of cross-frame attention [30, 43] and temporal window expansion [44] cannot reach the frames far from the query position, although the main action actually happens in the latter part of videos. The joint space-time attention model, on the other hand, is capable of global modeling but fails to focus on informative regions. In contrast, TC-CLIP consistently highlights the regions relevant to the query positions (e.g., hands and grabbed objects) throughout the video, leading to more accurate predictions.
Appendix 0.G Experimental Setup Details
0.G.1 Dataset Details
@@
0.G.2 Implementation Details
During the bipartite soft matching [16, 1], we start with the seed tokens arranged based on the [CLS] token attention values in each frame. These tokens are then divided into two sets by alternating positions. Subsequently, $r$ pairs of tokens with the highest cosine similarity are merged by averaging their features, and the remaining two sets are then concatenated back together. We set $r$ to 100 in practice. This process is repeated iteratively, employing a constant $r$ scheduling for every iteration with an exception in the final iteration to ensure that the number of final context tokens becomes $k$.
During the training, we sample 16 frames to form a video clip. During the evaluation, two temporal clips with one spatial crop (2 $\times$ 1 view) per video are sampled to produce a prediction unless otherwise stated. The learnable prompts are initialized with the prompt “a photo of a” following [52, 18], and the weight of the VP module is initialized with the weight from the last layer of the CLIP text encoder. For training recipes, we follow [35] for zero-shot, few-shot, and fully-supervised settings and follow [11] for base-to-novel generalization. By default, we use the AdamW optimizer with momentum betas of (0.9, 0.98) and a weight decay of 0.001. The VP module’s initial learning rate is 10×10$\times$10 × larger than the base learning rate in each setting. Training configurations and evaluation metrics in each protocol are specified below.
@@ 0.G.2 Implementation Details
Few-shot action recognition. We adopt the $K$-shot training splits from [35] that randomly sampled $K=2,4,8,16$ videos from each class on HMDB-51, UCF-101, and SSv2. The models are evaluated using the first validation split of HMDB-51 and UCF-101 and the full validation split of SSv2. The base learning rate is set to $2\times 10^{-6}$ and is decayed to $2\times 10^{-8}$. The batch size is 64, and the total number of epochs is set to 50, starting with 5 linear warmup epochs.